OpenStudy (anonymous):

Use trigonometric identities to transform one side of the equation into the other. (0 < θ , π/2) (tan θ + cot θ )/ tan θ = csc^2 θ Please I need to know how to work this out step by step.

OpenStudy (anonymous):

\[\begin{align*}\frac{\tan\theta+\cot\theta}{\tan\theta}&=\frac{\tan\theta}{\tan\theta}+\frac{\cot\theta}{\tan \theta}\\\\ &=1+\cot\theta\cdot\frac{1}{\tan\theta}\\\\ &=1+\cot\theta\cdot\cot\theta\\\\ &=1+\cot^2\theta\\\\ &=\cdots \end{align*}\]

OpenStudy (anonymous):

I don't understand what you did to get csc^2 θ .

OpenStudy (anonymous):

Please, I really need to know.

OpenStudy (anonymous):

K i got it. thanks.